# Enhanced Smart Contract Reputability Analysis using Multimodal Data Fusion on Ethereum

### v1

@@ Abstract
The evaluation of smart contract reputability is essential to foster trust in decentralized ecosystems. However, existing methods that rely solely on static code analysis or transactional data, offer limited insight into evolving trustworthiness. We propose a multimodal data fusion framework that integrates static code features with transactional data to enhance reputability prediction. Our framework initially focuses on static code analysis, utilizing GAN-augmented opcode embeddings to address class imbalance, achieving 97.67% accuracy and a recall of 0.942 in detecting illicit contracts, surpassing traditional oversampling methods. This forms the crux of a reputability-centric fusion strategy, where combining static and transactional data improves recall by 7.25% over single-source models, demonstrating robust performance across validation sets. By providing a holistic view of smart contract behaviour, our approach enhances the model’s ability to assess reputability, identify fraudulent activities, and predict anomalous patterns. These capabilities contribute to more accurate reputability assessments, proactive risk mitigation, and enhanced blockchain security.
@@ I Introduction
Current reputability assessment methods predominantly rely on either static code analysis or transactional data. Static code analysis evaluates the contract’s source code or bytecode for vulnerabilities and design flaws using tools such as SmartCheck [8] and Slither [4]. Although effective in detecting inherent security issues, static analysis provides a limited view, omitting insights from real-world usage and external factors. In contrast, transactional data analysis tracks runtime behaviours such as transaction volume and user activity, providing contextual insights into contract interactions [9, 10]. However, these approaches are reactive, only detecting issues after they occur, and lacks the interpretability needed to understand the underlying causes. These isolated approaches fail to capture the full spectrum of reputability, which encompasses both intrinsic code quality and dynamic behavioural patterns influenced by user interactions and external events [11, 12].
This work addresses this gap by proposing a multimodal data fusion framework that integrates static code features with dynamic transactional data to provide a comprehensive and evolving assessment of smart contract reputability. This research is guided by the following research question: Can a multimodal data fusion framework predict the evolution of smart contract reputability by integrating static code analysis with transactional data?
Our framework combines static opcode embeddings, augmented using Generative Adversarial Networks (GANs) to address class imbalance, with transactional data for anomaly detection. We propose the use of boosting algorithms [13], such as LightGBM, CatBoost, and XGBoost, to predict reputability based on static code analysis, achieving robust performance compared to traditional oversampling methods. Additionally, we use convolutional autoencoder-based models to fuse static and dynamic data, capturing the interplay between a contract’s internal logic and its interactions [6, 14]. This combined approach enhances anomaly detection and captures reputability shifts that single-source models often miss. The main contributions from this work are as follows:
  • A novel multimodal data fusion framework that integrates static opcode embeddings and transactional features for comprehensive smart contract reputability analysis, incorporating anomaly detection techniques to identify reputability shifts and improve robustness.
@@
By addressing the limitations of static-only and transactional-only approaches, this research lays the groundwork for further advancements in smart contract analysis. Our framework supports the proactive identification of reputability shifts over time and anomalous behaviour, contributing to the security and reliability of blockchain ecosystems.
The remainder of this paper is organized as follows: Section II reviews related work on smart contract reputability assessment and the limitations of static and transactional data approaches. Section III presents our multimodal data fusion framework, detailing the preprocessing steps, boosting algorithms, and fusion architecture. Section IV describes the experimental setup, datasets, and evaluation metrics, along with key performance comparisons. Section V discusses the limitations of our approach and future research directions. Finally, Section VI concludes the paper by summarizing our findings and emphasising the implications for blockchain security and trustworthiness.
@@ II-A Smart Contract Reputability
Traditional reputability assessments often focus on static code analysis, using tools like Slither [19] to identify coding errors and vulnerabilities. However, these methods are limited in scope, as they fail to account for runtime behaviours and interactions within a decentralized ecosystem. For example, a contract that passes initial security audits may later exhibit anomalous behaviour due to unexpected external interactions or changing usage patterns. Conversely, contracts perceived as risky at launch may build reputability over time through consistent, secure operation. This dynamic nature of smart contract reputability necessitates continuous monitoring that combines static analysis with insights from transactional data to capture changes in behaviour and performance [5].
Given the rise of DeFi exploits and their impact on user trust and market stability, it is imperative to move beyond isolated reputability assessments and adopt holistic frameworks that integrate static and dynamic data sources. Such an approach is essential not only for safeguarding assets but also for fostering a resilient decentralized ecosystem where participants can engage confidently with smart contract-based applications.
II-B Limitations of Static and Transactional Approaches
Smart contract reputability assessments typically rely on static code analysis or transactional data alone, yet both approaches have limitations that hinder a comprehensive evaluation of trustworthiness. Static code analysis tools, such as SmartCheck [8] and Slither [19], detect vulnerabilities in the source code pre-deployment, identifying issues like reentrancy flaws, unchecked calls, and arithmetic overflows. However, while effective at uncovering internal flaws in a contract’s logic, static analysis cannot capture external interactions or runtime behaviours, limiting its applicability in dynamic environments. Moreover, these tools rely on predefined vulnerability patterns, making them less effective at identifying novel attack vectors [20].
Static code analysis cannot account for runtime behaviours or unexpected interactions with external contracts. This limitation means that vulnerabilities arising from dynamic behaviours, such as gas usage anomalies or transaction sequencing issues, can go undetected, potentially affecting the reputability of the smart contract if exploited [21]. Moreover, static analysis is also constrained by predefined vulnerability patterns, making it less effective at identifying novel attack strategies. Recent studies [22] have shown that machine learning techniques, such as Long Short-term Memory (LSTMs) [23] and Extreme Gradient Boosting (XGBoost) [24], outperform symbolic tools by detecting complex vulnerabilities through pattern recognition in opcode sequences. However, these methods still fall short when contextual runtime data is unavailable.
@@ II-C Multimodal Learning for Anomaly Detection
Multimodal data fusion integrates information from multiple data sources to create a more comprehensive and accurate understanding of a phenomenon [26]. In the context of smart contracts, this involves combining static code features, such as opcode embeddings and vulnerability metrics, with dynamic transactional data to capture both pre-deployment vulnerabilities and runtime behaviours. By leveraging the strengths of each data type, multimodal learning provides richer insights that single-source methods cannot, revealing patterns and relationships that may be obscured when modalities are analysed in isolation [27, 28].
@@
The need for multimodal fusion becomes especially apparent when considering the limitations of static and transactional data-based approaches. Static code analysis identifies potential vulnerabilities in the contract’s internal logic but fails to capture dynamic behaviours that manifest only during runtime interactions. Conversely, transactional anomaly detection captures real-time behaviours but may flag anomalous patterns without indicating whether they stem from code-level flaws or external factors. By integrating static and dynamic features, multimodal approaches bridge this gap, providing context for behavioural anomalies and enabling more accurate and interpretable assessments.
@@
Multimodal fusion not only improves robustness by capturing complementary information but also reduces false positives by providing a holistic view of contract behaviour. By combining static and transactional data, it becomes possible to differentiate between legitimate spikes in transaction volume (e.g., protocol upgrades) and anomalies indicative of malicious behaviour. This integration supports proactive reputability assessments, addressing key challenges in maintaining trust and security within decentralized ecosystems.
@@ III-B Preprocessing
To prepare the dataset for analysis, several preprocessing steps were applied to both static code and transactional data. First, the bytecode sequences of smart contracts were disassembled into opcodes using the pyevmasm library. [^5] https://github.com/crytic/pyevmasm To streamline the analysis and reduce complexity, each opcode was categorised to a high-level category, such as arithmetic operations, logical operations, and memory operations, thereby preserving semantic information while simplifying the sequence structure.
The simplified opcode sequences were then converted into dense vector representations using a learned embedding approach. Each opcode was represented by a 50-dimensional vector through an embedding layer implemented in TensorFlow. By averaging these embeddings across the sequence, a single feature vector was generated for each smart contract, capturing nuanced relationships in the opcode sequence and facilitating robust static code analysis.
@@
III-C Boosted Static Code Analysis
To assess smart contract reputability using static code features, gradient boosting algorithms were employed due to their effectiveness in handling high-dimensional data. The training features included 50-dimensional opcode embeddings derived from simplified opcode sequences. These embeddings transformed categorical opcode data into dense vector representations, effectively capturing nuanced structural patterns within contracts.
@@ III-D Multimodal Data Fusion
Once the static code-level analysis was conducted, we transitioned to temporal analysis to examine how reputability evolves over time. Our focus then shifted to the characteristics of the transactions themselves. Consistent with the findings of [6] and [5], we concentrated primarily on normal and internal transactions since they were identified as particularly important for providing meaningful insights into the behaviour and reputability of the contracts.
We implemented a multimodal data fusion approach that integrates static code features with transactional data. The static features include 50-dimensional opcode embeddings derived from disassembled and simplified bytecode sequences, which capture the intrinsic structure and vulnerability-related characteristics of the smart contract. The transactional data comprises 820,920 transactions from reputable smart contracts (post-outlier removal) and 6,761 transactions from illicit ones, aggregated at an hourly granularity to balance temporal detail with computational feasibility.
@@
By incorporating opcode embeddings alongside transactional features, the CAE captures both runtime anomalies and underlying structural issues. This integration improves detection robustness, distinguishing between benign transaction spikes and illicit activities. The multimodal approach enables a holistic analysis of smart contract reputability, addressing limitations of static-only or transactional-only methods and providing richer insights into potential vulnerabilities and behavioural anomalies.
@@ IV Experimentation and Results
IV-A Static Code Analysis for Reputability Prediction
This section outlines the evaluation of boosting algorithms and data augmentation techniques for predicting smart contract reputability using static code features.
@@ IV-A Static Code Analysis for Reputability Prediction
Therefore, the results confirm that static code features, such as opcode embeddings, provide significant insights into smart contract reputability, enabling accurate detection of illicit contracts. However, despite the improvements, static code features alone may not fully capture behavioural shifts indicative of evolving reputability. This motivates the need for multimodal data fusion that integrates transactional data to enhance predictions, even in the absence of explicit reputability labels at specific timestamps.
@@ IV-B Multimodal Anomaly Detection
To address these limitations, we developed a multimodal CAE that integrates opcode embeddings with transactional features. This multimodal approach incorporates static and dynamic data, enabling the model to capture both structural vulnerabilities and behavioural anomalies. As illustrated in Figure 2, the reconstruction error distribution for reputable contracts narrowed, with the median error for reputable contracts decreasing significantly compared to the transaction-only model. This indicates that the inclusion of opcode embeddings enhanced the model’s ability to capture the nuances of normal smart contract behaviour. In contrast, reconstruction errors for illicit contracts remained notably high, emphasizing their anomalous nature and reinforcing the importance of opcode embeddings in detecting complex irregularities within smart contract behaviours.
@@
The performance comparison across different thresholds for the transaction-only and multimodal CAEs is summarized in Table V. At the optimal threshold of 90, the multimodal CAE improved recall for illicit contracts by 7.25%, demonstrating the benefits of integrating transactional data with static code features. This threshold was identified as optimal because it achieved the best trade-off between precision and recall, minimizing False Positives (FP) which are illicit contracts incorrectly classified as reputable—while maintaining high detection rates for illicit contracts.
This enhancement reduced FP and allowed the model to better distinguish true anomalies from benign deviations. As shown in Table V, the inclusion of opcode embeddings provided richer contextual information, highlighting the effectiveness of multimodal fusion for robust anomaly detection, particularly in scenarios where static features alone are insufficient for capturing reputability shifts.
@@
This enhanced separation in the latent space underscores the contribution of opcode embeddings in reinforcing class separability. By leveraging both static and dynamic features, the multimodal CAE learns richer latent representations that effectively capture behavioural and structural nuances in smart contract transactions.
In summary, the multimodal CAE outperforms the transaction-only CAE across all key metrics, achieving a recall of 0.974 and an F1-score of 0.980 for illicit contracts at the optimal threshold of 90. These results demonstrate the significant advantage of integrating opcode embeddings with transactional data for anomaly detection. The findings support the hypothesis that multimodal data fusion enhances smart contract reputability prediction by capturing both structural vulnerabilities and behavioural anomalies, demonstrating the value of combining static code analysis with dynamic transactional data for a comprehensive reputability assessment.
@@ V Limitations and Future Directions
The complexity of multimodal fusion also posed challenges. Integrating static and transactional features increased the computational complexity of the model and required extensive hyperparameter tuning to balance the representation of both data modalities. Computational constraints limited the scope of tuning experiments, potentially impacting full optimization. Moreover, the resource demands of the model may hinder real-time analysis across large blockchain networks. Addressing these challenges may involve developing more efficient fusion frameworks, such as lightweight architectures or distributed processing methods, to enhance scalability while preserving the benefits of multimodal integration.

### v2

@@ Abstract
The evaluation of smart contract reputability is essential to foster trust in decentralized ecosystems. However, existing methods that rely solely on analysis of code or transactional data, offer limited insight into evolving trustworthiness. We propose a multimodal data fusion framework that integrates code feature analysis with transactional data to enhance reputability prediction. Our framework initially focuses on AI-based code analysis, utilizing GAN-augmented opcode embeddings to address class imbalance, achieving 97.67% accuracy and a recall of 0.942 in detecting illicit contracts, surpassing traditional oversampling methods. This forms the crux of a reputability-centric fusion strategy, where combining code and transactional data improves recall by 7.25% over single-source models, demonstrating robust performance across validation sets. By providing a holistic view of smart contract behaviour, our approach enhances the model’s ability to assess reputability, identify fraudulent activities, and predict anomalous patterns. These capabilities contribute to more accurate reputability assessments, proactive risk mitigation, and enhanced blockchain security.
@@ I Introduction
Current reputability assessment methods predominantly rely on either code analysis or transactional data. Code analysis evaluates the contract’s source code or bytecode for vulnerabilities and design flaws using tools such as SmartCheck [8] and Slither [4]. Although effective in detecting inherent security issues, code analysis provides a limited view, omitting insights from real-world usage and external factors. In contrast, transactional data analysis tracks runtime behaviours such as transaction volume and user activity, providing contextual insights into contract interactions [9, 10]. However, these approaches are reactive, only detecting issues after they occur, and lacks the interpretability needed to understand the underlying causes. These isolated approaches fail to capture the full spectrum of reputability, which encompasses both intrinsic code quality and dynamic behavioural patterns influenced by user interactions and external events [11, 12].
This work addresses this gap by proposing a multimodal data fusion framework that integrates code features with dynamic transactional data to provide a comprehensive and evolving assessment of smart contract reputability. This research is guided by the following research question: Can a multimodal data fusion framework predict the evolution of smart contract reputability by integrating AI-based code analysis with transactional data?
Our framework combines opcode embeddings, augmented using Generative Adversarial Networks (GANs) to address class imbalance, with transactional data for anomaly detection. We propose the use of boosting algorithms [13], such as LightGBM, CatBoost, and XGBoost, to predict reputability based on code analysis, achieving robust performance compared to traditional oversampling methods. Additionally, we use convolutional autoencoder-based models to fuse code and dynamic data, capturing the interplay between a contract’s internal logic and its interactions [6, 14]. This combined approach enhances anomaly detection and captures reputability shifts that single-source models often miss. The main contributions from this work are as follows:
  • A novel multimodal data fusion framework that integrates opcode embeddings and transactional features for comprehensive smart contract reputability analysis, incorporating anomaly detection techniques to identify reputability shifts and improve robustness.
@@
By addressing the limitations of code analysis-only and transactional-only approaches, this research lays the groundwork for further advancements in smart contract analysis. Our framework supports the proactive identification of reputability shifts over time and anomalous behaviour, contributing to the security and reliability of blockchain ecosystems.
The remainder of this paper is organized as follows: Section II reviews related work on smart contract reputability assessment and the limitations of code analysis and transactional data approaches. Section III presents our multimodal data fusion framework, detailing the preprocessing steps, boosting algorithms, and fusion architecture. Section IV describes the experimental setup, datasets, and evaluation metrics, along with key performance comparisons. Section V discusses the limitations of our approach and future research directions. Finally, Section VI concludes the paper by summarizing our findings and emphasising the implications for blockchain security and trustworthiness.
@@ II-A Smart Contract Reputability
Traditional reputability assessments often focus on code analysis, using tools like Slither [19] to identify coding errors and vulnerabilities. However, these methods are limited in scope, as they fail to account for runtime behaviours and interactions within a decentralized ecosystem. For example, a contract that passes initial security audits may later exhibit anomalous behaviour due to unexpected external interactions or changing usage patterns. Conversely, contracts perceived as risky at launch may build reputability over time through consistent, secure operation. This dynamic nature of smart contract reputability necessitates continuous monitoring that combines code analysis with insights from transactional data to capture changes in behaviour and performance [5].
Given the rise of DeFi exploits and their impact on user trust and market stability, it is imperative to move beyond isolated reputability assessments and adopt holistic frameworks that integrate code analysis and dynamic data sources. Such an approach is essential not only for safeguarding assets but also for fostering a resilient decentralized ecosystem where participants can engage confidently with smart contract-based applications.
II-B Limitations of Code analysis-based and Transactional Approaches
Smart contract reputability assessments typically rely on code analysis or transactional data alone, yet both approaches have limitations that hinder a comprehensive evaluation of trustworthiness. Static code analysis tools, such as SmartCheck [8] and Slither [19], detect vulnerabilities in the source code pre-deployment, identifying issues like reentrancy flaws, unchecked calls, and arithmetic overflows. However, while effective at uncovering internal flaws in a contract’s logic, static analysis typically cannot capture external interactions or runtime behaviours (unless some form of runtime analysis or verification is used [20]), limiting its applicability in dynamic environments. Moreover, these tools often rely on predefined vulnerability patterns, making them less effective at identifying novel attack vectors [21].
Code analysis often cannot account for runtime behaviours or unexpected interactions with external contracts. This limitation means that vulnerabilities arising from dynamic behaviours, such as gas usage anomalies or transaction sequencing issues, can often go undetected, potentially affecting the reputability of the smart contract if exploited [22]. Moreover, analysis is also often constrained by predefined vulnerability patterns, making it less effective at identifying novel attack strategies. Recent studies [23] have shown that machine learning techniques, such as Long Short-term Memory (LSTMs) [24] and Extreme Gradient Boosting (XGBoost) [25], outperform symbolic tools by detecting complex vulnerabilities through pattern recognition in opcode sequences. However, these methods still fall short when contextual runtime data is unavailable.
@@ II-C Multimodal Learning for Anomaly Detection
Multimodal data fusion integrates information from multiple data sources to create a more comprehensive and accurate understanding of a phenomenon [27]. In the context of smart contracts, this involves combining code features, such as opcode embeddings and vulnerability metrics, with dynamic transactional data to capture both pre-deployment vulnerabilities and runtime behaviours. By leveraging the strengths of each data type, multimodal learning provides richer insights that single-source methods cannot, revealing patterns and relationships that may be obscured when modalities are analysed in isolation [28, 29].
@@
The need for multimodal fusion becomes especially apparent when considering the limitations of code and transactional data-based approaches. Code analysis identifies potential vulnerabilities in the contract’s internal logic but often fails to capture dynamic behaviours that manifest only during runtime interactions. Conversely, transactional anomaly detection captures real-time behaviours but may flag anomalous patterns without indicating whether they stem from code-level flaws or external factors. By integrating code and dynamic features, multimodal approaches bridge this gap, providing context for behavioural anomalies and enabling more accurate and interpretable assessments.
@@
Multimodal fusion not only improves robustness by capturing complementary information but also reduces false positives by providing a holistic view of contract behaviour. By combining code and transactional data, it becomes possible to differentiate between legitimate spikes in transaction volume (e.g., protocol upgrades) and anomalies indicative of malicious behaviour. This integration supports proactive reputability assessments, addressing key challenges in maintaining trust and security within decentralized ecosystems.
@@ III-B Preprocessing
To prepare the dataset for analysis, several preprocessing steps were applied to both code and transactional data. First, the bytecode sequences of smart contracts were disassembled into opcodes using the pyevmasm library. [^5] https://github.com/crytic/pyevmasm To streamline the analysis and reduce complexity, each opcode was categorised to a high-level category, such as arithmetic operations, logical operations, and memory operations, thereby preserving semantic information while simplifying the sequence structure.
The simplified opcode sequences were then converted into dense vector representations using a learned embedding approach. Each opcode was represented by a 50-dimensional vector through an embedding layer implemented in TensorFlow. By averaging these embeddings across the sequence, a single feature vector was generated for each smart contract, capturing nuanced relationships in the opcode sequence and facilitating code analysis.
@@
III-C Boosted Code Analysis
To assess smart contract reputability using code features, gradient boosting algorithms were employed due to their effectiveness in handling high-dimensional data. The training features included 50-dimensional opcode embeddings derived from simplified opcode sequences. These embeddings transformed categorical opcode data into dense vector representations, effectively capturing nuanced structural patterns within contracts.
@@ III-D Multimodal Data Fusion
Once the code-level analysis was conducted, we transitioned to temporal analysis to examine how reputability evolves over time. Our focus then shifted to the characteristics of the transactions themselves. Consistent with the findings of [6] and [5], we concentrated primarily on normal and internal transactions since they were identified as particularly important for providing meaningful insights into the behaviour and reputability of the contracts.
We implemented a multimodal data fusion approach that integrates code features with transactional data. The code features include 50-dimensional opcode embeddings derived from disassembled and simplified bytecode sequences, which capture the intrinsic structure and vulnerability-related characteristics of the smart contract. The transactional data comprises 820,920 transactions from reputable smart contracts (post-outlier removal) and 6,761 transactions from illicit ones, aggregated at an hourly granularity to balance temporal detail with computational feasibility.
@@
By incorporating opcode embeddings alongside transactional features, the CAE captures both runtime anomalies and underlying structural issues. This integration improves detection robustness, distinguishing between benign transaction spikes and illicit activities. The multimodal approach enables a holistic analysis of smart contract reputability, addressing limitations of code-only or transactional-only methods and providing richer insights into potential vulnerabilities and behavioural anomalies.
@@ IV Experimentation and Results
IV-A AI-based Code Analysis for Reputability Prediction
This section outlines the evaluation of boosting algorithms and data augmentation techniques for predicting smart contract reputability using code features.
@@ IV-A AI-based Code Analysis for Reputability Prediction
Therefore, the results confirm that code features, such as opcode embeddings, provide significant insights into smart contract reputability, enabling accurate detection of illicit contracts. However, despite the improvements, code features alone may not fully capture behavioural shifts indicative of evolving reputability. This motivates the need for multimodal data fusion that integrates transactional data to enhance predictions, even in the absence of explicit reputability labels at specific timestamps.
@@ IV-B Multimodal Anomaly Detection
To address these limitations, we developed a multimodal CAE that integrates opcode embeddings with transactional features. This multimodal approach incorporates code and dynamic data, enabling the model to capture both structural vulnerabilities and behavioural anomalies. As illustrated in Figure 2, the reconstruction error distribution for reputable contracts narrowed, with the median error for reputable contracts decreasing significantly compared to the transaction-only model. This indicates that the inclusion of opcode embeddings enhanced the model’s ability to capture the nuances of normal smart contract behaviour. In contrast, reconstruction errors for illicit contracts remained notably high, emphasizing their anomalous nature and reinforcing the importance of opcode embeddings in detecting complex irregularities within smart contract behaviours.
@@
The performance comparison across different thresholds for the transaction-only and multimodal CAEs is summarized in Table V. At the optimal threshold of 90, the multimodal CAE improved recall for illicit contracts by 7.25%, demonstrating the benefits of integrating transactional data with code features. This threshold was identified as optimal because it achieved the best trade-off between precision and recall, minimizing False Positives (FP) which are illicit contracts incorrectly classified as reputable—while maintaining high detection rates for illicit contracts.
This enhancement reduced FP and allowed the model to better distinguish true anomalies from benign deviations. As shown in Table V, the inclusion of opcode embeddings provided richer contextual information, highlighting the effectiveness of multimodal fusion for robust anomaly detection, particularly in scenarios where code features alone are insufficient for capturing reputability shifts.
@@
This enhanced separation in the latent space underscores the contribution of opcode embeddings in reinforcing class separability. By leveraging both code and dynamic features, the multimodal CAE learns richer latent representations that effectively capture behavioural and structural nuances in smart contract transactions.
In summary, the multimodal CAE outperforms the transaction-only CAE across all key metrics, achieving a recall of 0.974 and an F1-score of 0.980 for illicit contracts at the optimal threshold of 90. These results demonstrate the significant advantage of integrating opcode embeddings with transactional data for anomaly detection. The findings support the hypothesis that multimodal data fusion enhances smart contract reputability prediction by capturing both structural vulnerabilities and behavioural anomalies, demonstrating the value of combining AI-based code analysis with dynamic transactional data for a comprehensive reputability assessment.
@@ V Limitations and Future Directions
The complexity of multimodal fusion also posed challenges. Integrating code and transactional features increased the computational complexity of the model and required extensive hyperparameter tuning to balance the representation of both data modalities. Computational constraints limited the scope of tuning experiments, potentially impacting full optimization. Moreover, the resource demands of the model may hinder real-time analysis across large blockchain networks. Addressing these challenges may involve developing more efficient fusion frameworks, such as lightweight architectures or distributed processing methods, to enhance scalability while preserving the benefits of multimodal integration.